OpenStudy (cutiecomittee123):

for the value tan x/2 =1 which value is a solution pi/2 3pi/2 3pi/4 pi/4

jimthompson5910 (jim_thompson5910):

how far did you get here?

OpenStudy (cutiecomittee123):

Well I tried plugging in each of the values given into the x of the equation which failed miserably. I know that in terms of the unit circle we are looking for where tan is equal to one....so there are 4 points where that works but none of them are the values that are given.

jimthompson5910 (jim_thompson5910):

did you make sure your calculator was in radian mode?

OpenStudy (cutiecomittee123):

Well I used mathway, not my calculator. Which is actually a good idea let me try that.

OpenStudy (cutiecomittee123):

omg the answer is a. obviousely because pi/2 divided by 2 is 1. Well thats what I get when I plug it into the calculator.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (cutiecomittee123):

okay can you help me with another?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (cutiecomittee123):

evaluate Cos^-1 1/2

jimthompson5910 (jim_thompson5910):

ok use the unit circle and locate the angle that corresponds to the point with an x coordinate of x = 1/2

OpenStudy (cutiecomittee123):

Oh yeah okay. So its not the angle that is x=1/2 but it is the one that cooresponds with it?

jimthompson5910 (jim_thompson5910):

yeah you're basically asking yourself "what is theta when cos(theta) = 1/2 ?"

OpenStudy (cutiecomittee123):

Okay well let me take a look at the unit circle

OpenStudy (cutiecomittee123):

wouldnt it be 7pi/6

jimthompson5910 (jim_thompson5910):

look in Q1

OpenStudy (cutiecomittee123):

so then just pi/3?

OpenStudy (cutiecomittee123):

I was thinking it is the one opossite to that

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (cutiecomittee123):

okay hold on so the answer has to be in radians

OpenStudy (cutiecomittee123):

is pi/3 in radians??

jimthompson5910 (jim_thompson5910):

yes